# CryoSplat: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction

### v2

@@ Introduction
$\tilde{\mathcal{G}}(\tilde{\mathbf{r}}|\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}})=\int_{\mathbb{R}}\dot{\mathcal{G}}(\mathbf{r}|\dot{\boldsymbol{\mu}},\dot{\boldsymbol{\Sigma}})dr_{z}.$ (9)
@@
$),\tilde{\mathcal{G}}(\tilde{\mathbf{r}}|\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}})=\frac{1}{2\pi|\tilde{\boldsymbol{\Sigma}}|^{\frac{1}{2}}}\exp(-\frac{1}{2}(\tilde{\mathbf{r}}-\tilde{\boldsymbol{\mu}})^{\top}\tilde{\boldsymbol{\Sigma}}^{-1}(\tilde{\mathbf{r}}-\tilde{\boldsymbol{\mu}}$)), (10)
where $\tilde{\mathbf{r}}=[r_{x},r_{y}]^{\top}$ denotes the 2D Cartesian coordinates in real space. As discussed in (Zwicker et al. 2002), this projection can be computed analytically by removing the z-axis components from the mean and covariance:
$\begin{cases}\dot{\boldsymbol{\mu}}=[\dot{\mu}_{x},\dot{\mu}_{y},\dot{\mu}_{z}]^{\top}\Rightarrow[\dot{\mu}_{x},\dot{\mu}_{y}]^{\top}=\tilde{\boldsymbol{\mu}}\\ \dot{\boldsymbol{\Sigma}}=\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}&\dot{\sigma}_{xz}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}&\dot{\sigma}_{yz}\\ \dot{\sigma}_{xz}&\dot{\sigma}_{yz}&\dot{\sigma}_{zz}\\ \end{bmatrix}\Rightarrow\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}\\ \end{bmatrix}=\tilde{\boldsymbol{\Sigma}}\end{cases}$ (11)
In prior works, such as 3DGS, the normalization term $1/(2\pi|\tilde{\boldsymbol{\Sigma}}|^{\frac{1}{2}})$ is often omitted, as their primary focus is on photorealistic novel view synthesis rather than the physical fidelity of the underlying 3D representation. However, in cryo-EM reconstruction, the ultimate goal is to recover the correct 3D volume. Omitting this normalization introduces bias in amplitude and leads to incorrect reconstructions. Therefore, unlike 3DGS, we retain the normalization term to preserve the quantitative correctness of the model.
After projection, the final image is constructed by summing the weighted contributions of all projected 2D Gaussians and applying the PSF:
@@
In practice, consistent with typical cryo-EM pipelines, we apply the CTF instead in the Fourier domain:
@@
where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the Fourier and inverse Fourier transform operators, respectively.
Fast differentiable rasterization.
We adopt the efficient tile-based rasterization framework from 3DGS (Kerbl et al. 2023), which enables scalable and differentiable processing of tens of thousands of Gaussians via per-tile accumulation. Unlike 3DGS, which uses alpha blending for photorealistic rendering, we modify the rasterization to directly sum projected Gaussian contributions, in accordance with the physical transmission model in cryo-EM.
A key reason we choose to operate in real space is that it allows us to fully exploit this fast rasterization strategy. In high-resolution reconstructions, individual Gaussians in real space have small spatial scales and affect only a few nearby pixels. This locality means each GPU thread—responsible for a single pixel—only needs to process a small subset of all Gaussians. In contrast, Gaussians in Fourier space become broad as resolution increases, leading to near-global support. As a result, each pixel in the frequency domain must aggregate contributions from nearly all Gaussians, making fast rendering impractical.
For an image of shape $D\times D$, the original 3DGS implementation places the continuous coordinate center at $(D-1)/2,(D-1)/2]^{\top}$, i.e., halfway between two discrete pixels. In contrast, FFT-based image formation assumes the origin is located at the integer grid point $\lfloor D/2\rfloor,\lfloor D/2\rfloor]^{\top}$. To ensure compatibility with FFT-based forward and backward modeling, we shift the rasterization coordinates by half a pixel so that the image center aligns with the FFT grid.
This alignment eliminates phase inconsistencies and enables accurate electron projection simulation, while preserving the computational efficiency of the 3DGS architecture.
Loss function.
Unlike previous GMM-based methods, which often rely on specially designed loss functions combined with complex regularization or constraints to ensure stable optimization, we adopt a much simpler formulation. Specifically, we directly apply the mean squared error (MSE) loss between the GMM-based projection $\mathcal{I}$ and the observed image $\mathcal{I}_{\rm ref}$:
$\mathcal{L}=\|\mathcal{I}-\mathcal{I}_{\rm ref}\|_{2}.$ (14)
Experiment
Figure 2: Qualitative and quantitative comparison of voxel-based, neural, and GMM-based representations. (Left) Final 3D reconstructions on four real datasets visualized with ChimeraX (Pettersen et al. 2021). (Right) FSC curves are plotted for quantitative evaluation. Gray dashed lines indicate the standard resolution thresholds of $0.5$ and $0.143$, reported in Angstroms (Å). CryoSplat consistently achieves higher resolution across all datasets.
Figure 3: Evaluation of reconstruction performance with varying numbers of Gaussians. Increasing the number improves both accuracy and robustness, with 10,00010{,}000 Gaussians generally sufficient for most datasets.
Figure 4: Runtime efficiency comparison across reconstruction methods at different resolutions ($D=192$ and $D=256$). Frame rates (FPS) are measured under increasing numbers of Gaussians (log-scaled).
Experimental Settings
Datasets.
We evaluate our method on four publicly available cryo-EM datasets from the Electron Microscopy Public Image Archive (EMPIAR) (Iudin et al. 2016): EMPIAR-10028 (Pf80S ribosome) (Wong et al. 2014), EMPIAR-10049 (RAG complex) (Ru et al. 2015), EMPIAR-10076 (E. coli LSU assembly) (Davis et al. 2016), and EMPIAR-10180 (pre-catalytic spliceosome) (Plaschka, Lin, and Nagai 2017). These datasets span a range of structural complexity and image quality, from rigid assemblies with high contrast to highly heterogeneous macromolecular machines. For each dataset, we use the provided particle images, consensus pose estimates, and CTF parameters. All reconstructions are performed under the homogeneous assumption.
Evaluation metrics.
Since ground truth volumes are unavailable for real datasets, we follow standard practice and assess reconstruction quality using the gold standard Fourier Shell Correlation (FSC) (Van Heel and Schatz 2005). Each dataset is split evenly into two halves, and the method is applied independently to each subset. The resulting reconstructions are compared in Fourier space by computing FSC as a function of spatial frequency, which quantifies their consistency across frequency shells. The resolution is defined as the spatial frequency at which the FSC curve drops below the $0.143$ threshold.
Implementation details.
For all experiments, particle images from EMPIAR-10028, 10076, and 10180 are downsampled to $256\times 256$, while EMPIAR-10049 is used at its original $192\times 192$ resolution. Published particle translations are applied to the observed images via phase shifting in Fourier space prior to reconstruction, rather than through the GMM viewing transform. The 3D volume domain is defined as $0.5,0.5]^{3}$, assuming each projection lies within $0.5,0.5]^{2}$ in the image plane. Gaussians are initialized with random means $\boldsymbol{\mu}\sim\mathcal{N}(\mathbf{0},0.075^{2}\mathbf{I})$, isotropic scales $s_{x}\!=\!s_{y}\!=\!s_{z}\!=\!0.0075$, identity quaternion $\mathbf{q}\!=\![1,0,0,0]^{T}$, and amplitude $A=1/(2N)$, where $N$ is the number of Gaussians. All parameters are trainable. We use Adam (Kingma and Ba 2014) with batch size $1$, learning rate $0.001$, and exponential decay ($\gamma=0.1$) at each epoch. All GMMs are trained for $5$ epochs. For the voxel-based backprojection baseline, we use cryoDRGN’s official implementation (Zhong et al. 2021a). For neural representation learning, we follow cryoDRGN’s default configuration: three 1,0241{,}024-node layers, trained for $50$ epochs. Since related methods (2021a; 2021b; 2022a; 2022b; 2025) share similar neural fields and differ mainly in pose estimation, we focus comparison on cryoDRGN. All experiments are run on a single NVIDIA GeForce RTX 3090.
Evaluation on Real Datasets
We evaluate the performance of different volume representations on real cryo-EM datasets under a homogeneous reconstruction setting. To ensure a fair comparison focused solely on the choice of volume representation, all methods reconstruct consensus maps using the same set of published particle poses, without performing pose estimation. Our evaluation focuses on two aspects: (i) the ability to reconstruct high-resolution consensus maps, and (ii) robustness to noise and imperfect pose assignments. We compare voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al. 2021a), and our proposed GMM-based method cryoSplat. Visualizations of the reconstructed volumes are shown in Fig. 2, and spatial resolution is quantified using gold-standard FSC curves. Each volume of cryoSplat is represented using 30,00030{,}000 Gaussians.
EMPIAR-10028.
The Pf80S ribosome is relatively easy to reconstruct due to its high-contrast images and structurally stable particles. All methods achieve high-resolution results ($3.80$ Å) and strong FSC agreement across the spectrum. cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoSplat outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
EMPIAR-10049.
The RAG complex poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows. CryoSplat significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å. Unlike the baselines, cryoSplat reconstructs the DNA elements and the NBD with minimal artifacts. Its FSC curve remains consistently above those of other methods across all spatial frequencies, highlighting its robustness to pose inaccuracies and structural variability.
EMPIAR-10076.
This LSU assembly dataset contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging. Both FSC analysis and visualization show that cryoSplat is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
EMPIAR-10180.
The spliceosome dataset features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging. The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoSplat is more robust to such motions and achieves a resolution of $4.26$ Å. FSC analysis further confirms that cryoSplat significantly outperforms the baselines across the frequency range.
CryoSplat consistently converges within just $5$ epochs, with FSC curves from the $4$th and $5$th epochs tightly overlapping, indicating stable optimization and improved generalization in consensus reconstruction.
Ablation Studies
Number of Gaussians.
Fig. 3 shows the FSC curves for cryoSplat with varying numbers of Gaussians. In general, increasing the number of Gaussians leads to improved FSC, as a denser GMM provides greater representational capacity. While cryoSplat performs well on most datasets, its relative performance varies due to differences in structural complexity and dataset-specific challenges. On EMPIAR-10028, cryoSplat reaches a resolution of $3.8$ Å under all settings. While configurations with fewer than 10,00010{,}000 Gaussians exhibit lower FSC values than cryoDRGN and backprojection across most frequencies, the curves intersect at the highest frequency, indicating comparable final resolution. For EMPIAR-10049, all cryoSplat settings significantly outperform both backprojection ($4.00$ Å) and cryoDRGN ($4.07$ Å), achieving a resolution of $2.49$ Å. Moreover, the FSC curves of all cryoSplat variants remain consistently above those of the two baselines across the entire frequency range. For EMPIAR-10076, the 30,00030{,}000-Gaussian model clearly outperforms other settings; even with fewer Gaussians, cryoSplat still surpasses the baselines, reaching $3.3$ Å. For EMPIAR-10180, the models with 10,00010{,}000 and 30,00030{,}000 Gaussians achieve the best FSC, reaching $4.3$ Å, while sparser GMMs remain competitive at high spatial frequencies. Overall, we find that using 10,00010{,}000 Gaussians is sufficient to reconstruct high-resolution volumes that consistently outperform baseline methods across most datasets.
Runtime efficiency.
[…]
Conclusion
@@
Appendix A Dataset Details

### v3

@@ 3.3.3 Orthogonal projection
$\tilde{G}(\tilde{{\bm{r}}}|\tilde{{\bm{\mu}}},\tilde{{\bm{\Sigma}}})=\int_{{\mathbb{R}}}\dot{G}({\bm{r}}|\dot{{\bm{\mu}}},\dot{{\bm{\Sigma}}})dr_{z}.$ (7)
@@
$),\tilde{G}(\tilde{{\bm{r}}}|\tilde{{\bm{\mu}}},\tilde{{\bm{\Sigma}}})=\frac{1}{2\pi|\tilde{{\bm{\Sigma}}}|^{\frac{1}{2}}}\exp\Big(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$$\Big)$, (8)
where $\tilde{{\bm{r}}}=[r_{x},r_{y}]^{\top}$ denotes the 2D Cartesian coordinates in real space.
In prior works, such as 3DGS, the normalization term $1/(2\pi|\tilde{{\bm{\Sigma}}}|^{\frac{1}{2}})$ is often omitted, as their primary focus is on photorealistic novel view synthesis rather than the physical fidelity of the underlying 3D representation. However, in cryo-EM reconstruction, the ultimate goal is to recover the correct 3D volume. Omitting this view-dependent normalization introduces bias in amplitude and leads to incorrect reconstructions. Therefore, unlike 3DGS, we retain the normalization term to preserve the quantitative correctness of the model.
After projection, the final image is constructed by summing the weighted contributions of all splats and applying the PSF:
@@
3.3.4 Fast differentiable rasterization
We adopt the efficient tile-based rasterization framework from Kerbl et al. (2023), which enables scalable and differentiable processing of tens of thousands of Gaussians via per-tile accumulation. Unlike 3DGS, which uses alpha blending for photorealistic rendering, we modify the rasterization to directly sum contributions of splats, in accordance with the physical transmission model in cryo-EM.
For an image ${\bm{X}}\!\in\!{\mathbb{R}}^{D\times D}$, the original 3DGS implementation places the continuous coordinate center at $(D-1)/2,(D-1)/2]^{\top}$, i.e., halfway between two discrete pixels. In contrast, FFT-based image formation assumes the origin is located at the integer grid point $\lfloor D/2\rfloor,\lfloor D/2\rfloor]^{\top}$. To ensure compatibility with FFT-based forward and backward modeling, we shift the rasterization coordinates by half a pixel so that the image center aligns with the FFT grid. This alignment eliminates phase inconsistencies and enables accurate electron projection simulation, while preserving the computational efficiency of the 3DGS architecture. Let ${\bm{X}},{\bm{Y}}\!\in\!\mathbb{R}^{D\times D}$ be the matrices representing the GMM-based projection $X$ and the observed image $Y$ after rasterization, respectively.
3.3.5 Loss function
Unlike previous GMM-based methods that rely on specially designed losses with complex regularization or constraints to ensure stable optimization, we adopt a much simpler formulation. Specifically, we directly apply the mean squared error (MSE) loss between the GMM-based projection $\bm{X}$} and the observed image $\bm{Y}$}: $\mathcal{L}=\|{\bm{X}}-{\bm{Y}}\|_{2}^{2}$.
4 Experiment
Figure 2: Qualitative and quantitative comparison of voxel-based, neural, and GMM-based representations. (Left) Final 3D reconstructions on four real datasets visualized with ChimeraX (Pettersen et al., 2021). (Right) FSC curves are plotted for quantitative evaluation. Gray dashed lines indicate the standard resolution thresholds of $0.5$ and $0.143$, reported in Angstroms (Å). CryoSplat consistently achieves higher resolution across all datasets.
Figure 3: Reconstruction performance with varying numbers of Gaussians. Increasing the number improves accuracy and robustness, with 10,00010{,}000 Gaussians generally sufficient for most datasets.
Figure 4: Runtime efficiency across reconstruction methods at different resolutions ($D=192$ and $D=256$). Frame rates (FPS) are measured under increasing numbers of Gaussians (log-scaled).
4.1 Experimental settings
Datasets. We evaluate our method on four publicly available cryo-EM datasets from the Electron Microscopy Public Image Archive (EMPIAR) (Iudin et al., 2016): EMPIAR-10028 (Pf80S ribosome) (Wong et al., 2014), EMPIAR-10049 (RAG complex) (Ru et al., 2015), EMPIAR-10076 (E. coli LSU assembly) (Davis et al., 2016), and EMPIAR-10180 (pre-catalytic spliceosome) (Plaschka et al., 2017). These datasets span a range of structural complexity and image quality, from rigid assemblies with high contrast to highly heterogeneous macromolecular machines. For each dataset, we use the provided particle images, consensus pose estimates, and CTF parameters. All reconstructions are performed under the homogeneous assumption.
Evaluation metrics. Since ground truth volumes are unavailable for real datasets, we follow standard practice and assess reconstruction quality using the gold standard Fourier Shell Correlation (FSC) (Van Heel & Schatz, 2005). Each dataset is split evenly into two halves, and the method is applied independently to each. Reconstructed volumes are compared in Fourier space by computing FSC as a function of spatial frequency, quantifying their consistency across frequency shells. Resolution is defined as the spatial frequency where the FSC curve drops below the $0.143$ threshold.
Implementation details. For all experiments, particle images from EMPIAR-10028, 10076, and 10180 are downsampled to $256\times 256$, while EMPIAR-10049 is used at its original $192\times 192$ resolution. Published particle translations are applied to the observed images via phase shifting in Fourier space prior to reconstruction, rather than through the GMM viewing transform. We do not apply any windowing to the observed particle images during preprocessing. The 3D volume is defined over the domain $E,E]^{3}$, and each 2D projection is assumed to lie within $E,E]^{2}$ in the image plane, where $E=0.5$ defines the spatial extent. Gaussians are initialized with random means ${\bm{\mu}}\sim\mathcal{N}({\bm{0}},0.075^{2}{\bm{I}})$, isotropic scales $s_{x}\!=\!s_{y}\!=\!s_{z}\!=\!0.0075$, identity quaternion ${\bm{q}}\!=\![1,0,0,0]^{T}$, and amplitude $A=1/(2N)$, where $N$ is the number of Gaussians. All parameters are trainable. We use Adam (Kingma & Ba, 2014) with batch size $1$, learning rate $0.001$, and exponential decay ($\gamma=0.1$) at each epoch. All GMMs are trained for $5$ epochs. For the voxel-based backprojection baseline, we use the implementation provided by Zhong et al. (2021a). For neural representation learning, we follow cryoDRGN’s default configuration: three 1,0241{,}024-node layers, trained for $50$ epochs. All experiments are run on a single NVIDIA GeForce RTX 3090.
4.2 Evaluation on real datasets
We evaluate the performance of different volume representations on real cryo-EM datasets under a homogeneous reconstruction setting. To ensure a fair comparison focused solely on the choice of volume representation, all methods reconstruct consensus maps using the same set of published particle poses, without performing pose estimation. Our evaluation focuses on two aspects: (i) the ability to reconstruct high-resolution consensus maps, and (ii) robustness to noise and imperfect pose assignments. Since related methods (Zhong et al., 2021a; b; Levy et al., 2022a; b; 2025) adopt cryoDRGN’s neural field implementation and differ mainly in pose estimation, we focus our comparison on cryoDRGN. Accordingly, we evaluate three representative approaches: voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al., 2021a), and our proposed GMM-based method cryoSplat. Visualizations of the reconstructed volumes are shown in Fig. 2, and spatial resolution is quantified using gold-standard FSC curves. Each volume of cryoSplat is represented using 30,00030{,}000 Gaussians.
The Pf80S ribosome (EMPIAR-10028) is relatively easy to reconstruct due to its high-contrast images and structurally stable particles. All methods achieve high-resolution results ($3.80$ Å) and strong FSC agreement across the spectrum. cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoSplat outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
The RAG complex (EMPIAR-10049) poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows. CryoSplat significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å. Unlike the baselines, cryoSplat reconstructs the DNA elements and the NBD with minimal artifacts. Its FSC curve remains consistently above those of other methods across all spatial frequencies, highlighting its robustness to pose inaccuracies and structural variability.
This LSU assembly dataset (EMPIAR-10076) contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging. Both FSC analysis and visualization show that cryoSplat is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
The spliceosome dataset (EMPIAR-10180) features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging. The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoSplat is more robust to such motions and achieves a resolution of $4.26$ Å. FSC analysis further confirms that cryoSplat significantly outperforms the baselines across the frequency range.
CryoSplat consistently converges within 5 epochs, with FSC curves from the 4th and 5th epochs tightly overlapping, indicating stable optimization and improved generalization.
4.3 Ablation studies
This section reports ablation studies of our approach. More results can be found in Appendix E.
Number of Gaussians. Fig. 3 shows the FSC curves for cryoSplat with varying numbers of Gaussians. In general, increasing the number of Gaussians leads to improved FSC, as a denser GMM provides greater representational capacity. While cryoSplat performs well on most datasets, its relative performance varies due to differences in structural complexity and dataset-specific challenges. On EMPIAR-10028, cryoSplat reaches a resolution of $3.8$ Å under all settings. While configurations with fewer than 10,00010{,}000 Gaussians exhibit lower FSC values than cryoDRGN and backprojection across most frequencies, the curves intersect at the highest frequency, indicating comparable final resolution. For EMPIAR-10049, all cryoSplat settings significantly outperform both backprojection ($4.00$ Å) and cryoDRGN ($4.07$ Å), achieving a resolution of $2.49$ Å. Moreover, the FSC curves of all cryoSplat variants remain consistently above those of the two baselines across the entire frequency range. For EMPIAR-10076, the 30,00030{,}000-Gaussian model clearly outperforms other settings; even with fewer Gaussians, cryoSplat still surpasses the baselines, reaching $3.3$ Å. For EMPIAR-10180, the models with 10,00010{,}000 and 30,00030{,}000 Gaussians achieve the best FSC, reaching $4.3$ Å, while sparser GMMs remain competitive at high spatial frequencies. Overall, we find that using 10,00010{,}000 Gaussians is sufficient to reconstruct high-resolution volumes that consistently outperform baseline methods across most datasets. Associated qualitative comparisons are provided in Appendix E.
@@ 4.3 Ablation studies
5 Conclusion
@@
Appendix A Details of method
A.1 Real space reconstruction
According to the Fourier slice theorem (Bracewell, 1956), illustrated in Fig. 1(c), the 2D Fourier transform of a projection corresponds to a central slice of the 3D Fourier transform of the volume, orthogonal to the projection direction and passing through the origin. Based on this property, an alternative and widely adopted formulation models reconstruction directly in the Fourier domain, where the image formation model becomes:
$\widehat{Y}(k_{x},k_{y})=\widehat{H}(k_{x},k_{y})\cdot\widehat{V}({\bm{W}}^{\top}{\bm{k}})\cdot e^{-2\pi i{\bm{k}}^{\top}{\bm{t}}}+\widehat{{\textnormal{$\epsilon$}}}$, (10)
where ${\bm{k}}=[k_{x},k_{y},0]^{\top}$ denotes the Cartesian coordinates in Fourier space, and the 2D spectrum $\widehat{Y}$, the CTF $\widehat{H}$ and the 3D spectrum $\widehat{V}$ denote the Fourier transform of $Y$, $H$ and $V$, respectively. The noise term $\widehat{{\textnormal{$\epsilon$}}}$ is similarly modeled as independent, zero-mean Gaussian noise in the Fourier domain.
A key reason we choose to operate in real space is that it allows us to fully exploit the fast rasterization strategy from 3DGS. In high-resolution reconstructions, individual Gaussians in real space have small spatial scales and affect only a few nearby tiles, as shown in Fig. 5(a). This locality means that each GPU thread is responsible for a single pixel and only needs to process a small subset of all Gaussians. In contrast, Gaussians in Fourier space become broad as resolution increases, leading to near-global support. As a result, each pixel in the frequency domain must aggregate contributions from nearly all Gaussians, making fast rendering impractical.
Figure 5: Rasterization details. (a) A Gaussian with small spatial scales in real space during high-resolution reconstruction overlaps at most four tiles, while in Fourier space it exhibits nearly global support. Tile boundaries are indicated by lines. (b) For a $4\times 4$ image, the origin of the continuous coordinate system during rasterization is defined differently: FFT-based coordinates place it at $2,2]^{\top}$, whereas image-centered coordinates place it at $1.5,1.5]^{\top}$. The origin is marked by a dot. Image-centered coordinates induce a phase error of $\pi{\bm{k}}/D$, up to $\pm\pi/2$ at the Nyquist frequency, degrading reconstruction at high frequency. The effect diminishes with larger $D$ but never vanishes.
A.2 Computational details
Section 3.3.3 describes how a 3D Gaussian is projected along the $z$-axis to form a splat. As discussed in Zwicker et al. (2002), the splat can be computed analytically by removing the $z$-axis components from the mean and covariance:
$\begin{cases}\dot{{\bm{\mu}}}=[\dot{\mu}_{x},\dot{\mu}_{y},\dot{\mu}_{z}]^{\top}\Rightarrow[\dot{\mu}_{x},\dot{\mu}_{y}]^{\top}=\tilde{{\bm{\mu}}}\\ \dot{{\bm{\Sigma}}}=\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}&\dot{\sigma}_{xz}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}&\dot{\sigma}_{yz}\\ \dot{\sigma}_{xz}&\dot{\sigma}_{yz}&\dot{\sigma}_{zz}\\ \end{bmatrix}\Rightarrow\begin{bmatrix}\dot{\sigma}_{xx}&\dot{\sigma}_{xy}\\ \dot{\sigma}_{xy}&\dot{\sigma}_{yy}\\ \end{bmatrix}=\tilde{{\bm{\Sigma}}}\end{cases}$ (11)
thereby enabling an efficient computation of Eq. (7).
As discussed in Sec. 3.3.4 and shown in Fig. 5(b), when a continuous image $X:{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ is rasterized onto pixels ${\bm{X}}\!\in\!{\mathbb{R}}^{D\times D}$, the origin of the continuous coordinate system should be aligned with $\lfloor D/2\rfloor,\lfloor D/2\rfloor]^{\top}$ to match the FFT-based coordinate convention used in cryo-EM. Formally,
${X}_{i,j}=X\left((j-\lfloor\frac{D}{2}\rfloor)\frac{2E}{D},-(i-\lfloor\frac{D}{2}\rfloor)\frac{2E}{D}\right),$ (12)
where ${X}_{i,j}$ denotes $(i,j)$-th entry of matrix $\bm{X}$}. Note that the row and column indices correspond to the $y$- and $x$-axes, respectively, with the $y$-axis flipped during this rasterization.
In practice, since the PSF corresponds to a large convolution kernel, we apply the contrast transfer function (CTF) in the Fourier domain after rasterization for efficiency:
${X}_{i,j}=\mathcal{F}^{-1}\left(\widehat{{\bm{H}}}\odot\mathcal{F}\left(\sum_{i=1}^{N}A_{i}\tilde{G}_{i}\left((j-\lfloor\tfrac{D}{2}\rfloor)\tfrac{2E}{D},-(i-\lfloor\tfrac{D}{2}\rfloor)\tfrac{2E}{D}\right)\right)\right),$ (13)
where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the Fourier and inverse Fourier transform operators, respectively. $\widehat{{\bm{H}}}$ is the rasterized CTF $\widehat{H}$ and $\odot$ denotes element-wise (Hadamard) product.
Before deriving the gradients, we first define
[…]
which is the pre-rasterization continuous image. For clarity, we omit the Gaussian index $i$ in the following derivations, as the gradients are computed independently for each Gaussian. We denote by $\mathbb{P}$} the set of 2D coordinates corresponding to the centers of rasterized pixels. When $\tilde{{\bm{r}}}\in{\mathbb{P}}$, the coordinate $\tilde{{\bm{r}}}=[r_{x},r_{y}]^{\top}$ refers to a discrete sampling location in the image plane. The gradients used in the backward pass can be summarized as
$\left\{\begin{aligned} \frac{\partial\mathcal{L}}{\partial A}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\frac{\partial Q(\tilde{{\bm{r}}})}{\partial A}\\ \nabla_{\bm{\mu}}\mathcal{L}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\nabla_{\bm{\mu}}Q(\tilde{{\bm{r}}})\\ \nabla_{\bm{s}}\mathcal{L}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\nabla_{\bm{\Sigma}}Q(\tilde{{\bm{r}}})\circ\frac{\partial{\bm{\Sigma}}}{\partial{\bm{s}}}\\ \nabla_{\bm{q}}\mathcal{L}&=\sum_{\tilde{{\bm{r}}}\in{\mathbb{P}}}\frac{\partial\mathcal{L}}{\partial Q(\tilde{{\bm{r}}})}\nabla_{\bm{\Sigma}}Q(\tilde{{\bm{r}}})\circ\frac{\partial{\bm{\Sigma}}}{\partial{\bm{q}}}\end{aligned}\right$. (15)
where $\circ$ denotes the composition of Jacobian operators (chain rule). The derivation of gradients with respect to the amplitude $A$ and mean $\bm{\mu}$} is trivial, which can be given directly by
[…]
and
$\left\{\begin{aligned} \nabla_{\tilde{{\bm{\mu}}}}Q&=A\tilde{G}(\tilde{{\bm{r}}})\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})\\ \nabla_{\bm{\mu}}Q&=W[\nabla_{\tilde{{\bm{\mu}}}}Q^{\top}\;0]^{\top}\end{aligned}\right$. (17)
[…]
For completeness, we provide the derivation of the covariance gradients $\nabla_{\bm{\Sigma}}Q$, noting that our formulation retains the normalization term, which is omitted in 3DGS (Kerbl et al., 2023). Remember
[…]
$).\displaystyle=\frac{|\tilde{{\bm{\Sigma}}}^{-1}|^{\frac{1}{2}}}{2\pi}\exp(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$)).
We can first compute
$\nabla_{\tilde{{\bm{\Sigma}}}^{-1}}Q$ $=A\exp(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}))\frac{1}{4\pi}|\tilde{{\bm{\Sigma}}}^{-1}|^{-\frac{1}{2}}|\tilde{{\bm{\Sigma}}}^{-1}|\tilde{{\bm{\Sigma}}}^{\top}$ (19)
$)\displaystyle+A\frac{|\tilde{{\bm{\Sigma}}}^{-1}|^{\frac{1}{2}}}{2\pi}\exp(-\frac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})^{\top}\tilde{{\bm{\Sigma}}}^{-1}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}))(-\tfrac{1}{2}(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$$^{\top})$
$),\displaystyle=\tfrac{1}{2}A\tilde{G}(\tilde{{\bm{r}}})(\tilde{{\bm{\Sigma}}}-(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}})(\tilde{{\bm{r}}}-\tilde{{\bm{\mu}}}$$^{\top})$,
and then $\nabla_{\tilde{{\bm{\Sigma}}}}Q=-\tilde{{\bm{\Sigma}}}^{-\top}\nabla_{\tilde{{\bm{\Sigma}}}^{-1}}Q\tilde{{\bm{\Sigma}}}^{-\top}$. Finally,
$\nabla_{\dot{{\bm{\Sigma}}}}Q=\begin{bmatrix}\nabla_{\tilde{{\bm{\Sigma}}}}Q&{\bm{0}}\\[6.0pt] {\bm{0}}^{\top}&0\end{bmatrix}$. (20)
The subsequent derivations of $\nabla_{\bm{s}}\mathcal{L}$ and $\nabla_{\bm{q}}\mathcal{L}$ follow exactly the formulation in Kerbl et al. (2023).
Appendix B Dataset details